OpenStudy (anonymous):

I need help indentifying the natural domain of a function, The function f(x) is defined as F(x) = 3x+4/5x+9 The question asks me, The domain of (f) contains all real numbers (X) except X = ???????

OpenStudy (anonymous):

except the numbers that make 5x+9=0 ,that is x=-9/5 because if 5x+9=0 then f(x) is undefined.

OpenStudy (anonymous):

ok thanks that makes sense!!